when the terminal potential difference of a real cell is less than its emf
1
Expert's answer
2012-06-01T08:43:55-0400
A source of emf also has some resistance which is due to electrolytes
and
electrodes. Therefore some useful energy of emf-source is used in
passing the
current through the source.
Let the internal resistance of the source is
“r”
Potential drop across emf source = Vr
Since V = E - Vr, so V = E -
Ir.
Where V = terminal potential difference
When no current is drawn from
the source then emf and terminal
potential difference are equal. If the
current is passing through the
circuit then the terminal potential difference
is less than that of
the emf of the source.